### v1

# XXLTraffic: Expanding and Extremely Long Traffic Dataset for Ultra-Dynamic Forecasting Challenges

Abstract
Traffic forecasting is crucial for smart cities and intelligent transportation initiatives, where deep learning has made significant progress in modeling complex spatio-temporal patterns in recent years. However, current public datasets have limitations in reflecting the ultra-dynamic nature of real-world scenarios, characterized by continuously evolving infrastructures, varying temporal distributions, and temporal gaps due to sensor downtimes or changes in traffic patterns. These limitations inevitably restrict the practical applicability of existing traffic forecasting datasets. To bridge this gap, we present XXLTraffic, the largest available public traffic dataset with the longest timespan and increasing number of sensor nodes over the multiple years observed in the data, curated to support research in ultra-dynamic forecasting. Our benchmark includes both typical time-series forecasting settings with hourly and daily aggregated data and novel configurations that introduce gaps and down-sample the training size to better simulate practical constraints. We anticipate the new XXLTraffic will provide a fresh perspective for the time-series and traffic forecasting communities. It would also offer a robust platform for developing and evaluating models designed to tackle ultra-dynamic and extremely long forecasting problems. Our dataset supplements existing spatio-temporal data resources and leads to new research directions in this domain.
1 Introduction
Traffic prediction plays a vital role in intelligent transportation systems by offering accurate and dependable forecasts of future traffic conditions that serve as critical references for route design, urban construction, and road management. The long-term spatio-temporal relationship is essential and irreplaceable for traffic forecasting, i.e., the long-term relationship can not only provide context that helps understand anomalies or irregularities in short-term data. but also help identify patterns and trends over different timescales, such as daily, weekly, monthly, and yearly cycles. In recent years, significant work has focused on short-term and long-term traffic flow prediction. Various deep learning techniques, such as GNNs, emphasize extracting spatial relationships in traffic networks [13], while Transformer variants and upgraded architectures and methods focus on capturing short-term and long-term temporal relationships [29]. Existing work has achieved promising performance in both short-term and long-term predictions, but this is limited to relatively simple experimental setups.
The inevitable tendency in intelligent transportation systems is to design models that incorporate more realistic circumstances rather than just simple experimental conditions. This requires the use of data sets that accurately reflect real-world scenarios. This serves as a strong incentive for us to create and develop XXLTraffic, i.e., why does XXLTraffic focus on ultra-dynamic forecasting? Why is it necessary to include extremely long-term data? We will introduce the commonalities and developments of existing datasets, as well as the difficulties in establishing XXLTraffic more specifically.
Figure 1: Our dataset are evolving and longer than existing datasets. Existing datasets are either limited by short temporal spans or insufficient spatial nodes. In contrast, our dataset features an evolving growth of spatial nodes and spans over 20 years.
Recent Advances in Expanding Traffic Datasets. Real-world traffic scenarios necessitate more complex prediction settings, involving extended temporal horizons or broader spatial coverage in experiments. In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  [30] and [11] expanded input and output lengths up to four times on existing datasets.From a spatial perspective, [4] published a dataset with nodes growing annually and provided an evolving network to support new node predictions. [33] proposed a continual learning framework with pattern expansion mechanisms based on [4]. Additionally, [27] and [25] offered larger-scale spatial node datasets to support subsequent researchers. Recent work has explored longer temporal step experimental settings and released traffic datasets spanning up to five years and thousands of nodes. However, in specific scenarios, such as future traffic prediction for highway planning, these data and experimental settings fall short. As shown in Figure 1, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting. Highway construction typically spans several years, requiring predictions several years in advance. These extremely long prediction scenarios necessitate traffic datasets with even longer temporal spans for support. As the temporal span extends, urban infrastructure development and road construction can lead to shifts in traffic patterns, resulting in an evolving domain shift. This observation motivated us to provide an expanding and extremely long traffic dataset. Additionally, the combination of these factors enables the extraction of more temporal patterns from extremely long sequences, allowing for the possibility of longer input sequences.
Challenges. The ultra-dynamic challenge encompasses three key aspects: (1) Continuously evolving states of the underlying spatio-temporal infrastructures, characterized by an expanding number of nodes over the years. This continuous growth introduces complexity as the infrastructure adapts and expands. (2) Evolving temporal distributions over an extremely long observation period, which is crucial for extremely long forecasting. This requires models to adapt to changes in patterns and trends over extensive temporal spans. (3) Ultra spatio-temporal dynamicity with temporal gaps. These gaps can occur due to various reasons such as sensor data becoming unavailable during road construction, maintenance activities, or shifts in data distributions. This dynamic nature necessitates robust modeling techniques to handle inconsistencies and interruptions in data.
Contribution. We have constructed a traffic dataset with an exceptionally long temporal span and broader regional coverage, providing aggregated data and benchmarking, as well as a benchmarking setup considering extremely long prediction scenarios for future exploration:
  • We propose XXLTraffic, a dataset that spans up to 23 years and exhibits evolutionary growth. It includes data from 9 regions, with detailed data collection and processing procedures for expansion and transformation. This dataset supports both temporally scalable and spatially scalable challenges in traffic prediction.
  • We present an experimental setup with temporal gaps for extremely long prediction with gaps scenarios and provide a benchmark of aggregated versions of hourly and daily datasets.
  • We provide the exploration of input features through evolving temporal distributions over an extremely long observation period. Additionally, our datasets support zero-shot forecasting for new sensors.
2 Preliminaries
In this section, we will define traffic data and traffic prediction tasks.
Definition 1. Traffic datasets:
Traffic data primarily consists of vehicle flow detection data collected by sensors distributed across various locations in the traffic network. It is generally represented by $X_{i}$in $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors, and $C$ denotes the number of features.
Definition 2. Short-term traffic prediction:
Short-term traffic prediction primarily focuses on forecasting traffic speed or flow within the next hour. As shown in Equation 1, the input length $\alpha$ and output length $\beta$ are generally set to 12 steps.
$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$ (1)
Definition 3. Long-term multivariate prediction:
This task mainly focuses on long-sequence time series prediction, which includes the traffic dataset. As shown Table 1, the sequence length can reach up to 2880 steps.
Definition 4. Extremely Long Prediction with Gaps:
Based on Equation 1, the observation and prediction are not adjacent but are instead separated by a gap period $g$, as shown in the following formula.
$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$ (2)
3 Gaps and Comparison with Existing Traffic Datasets
Figure 2: Our expanding and extremely long prediction addresses the existing limitations in both short-term and long-term predictions.
As shown in Table 1, existing traffic prediction work can easily be divided into short-term and long-term settings. The short-term setting originated from the STGCN[41] work, while the long-term setting was first introduced by LSTNet [14] and subsequently established as a widely adopted experimental framework by Informer [45]. In recent years, short-term prediction typically has a maximum step length of 12 steps, while long-term prediction reaches up to 720 steps. However, works such as Witran [11] and DAN [19] recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length. Despite the differences in step lengths, their observed and predicted values are concatenated tightly together, as shown in Equation 1. To accommodate complex real-world scenarios, such as highway route planning predictions, it is necessary to introduce a gap of several years between observation and prediction. Typically, existing datasets lack the temporal coverage required to support gaps exceeding one year.At the same time, predicting several years in advance also implies the need to forecast traffic for sensors at new locations, taking into account the evolving nature of the road network. Even when such coverage is available, works like [33] and [4] utilize evolving datasets but do not provide sufficient data to train models for extended durations. To overcome these gaps, our Expanding and Extremely Long Traffic Dataset robustly supports these complex scenarios.
Table 1: Summary of recent short-term traffic forecasting and long-term multivariate forecasting
Datasets Model Series Length
Short-term STGCN [41] {3,6,9,12}
DCRNN [20], GWN [40], BTF [3], DMSTGCN [10], {3,6,12}
GTS [28], STGODE [7], PM-MemNet [16], STAEFormer [23]
AGCRN [1], STSGCN [32], ,DSTAGNN [15] D2STGNN [31], {12}
DyHSL [44],PDFormer [12], MultiSPANS [47], GMSDR [22]
Long-term MTGNN [39],LSTNet [14] {3,6,12,24}
ARU [6] {12,24,48,168,336}
LogSparse_Trans [18] {24,48,72,96,120,144,168,192}
AST [38] {8,24,168,336}
SSDNet [21] {20,24,30,138}
Informer [45], Autoformer [37], FEDformer [46], {24,48,96,192,336,720}
Linear [17], Triformer [5], Pyraformer [24]
DSformer [42],DeepTime [35],DLinear [43]
Witran [11] {168, 336, 720, 1440, 2880}
DAN [19] {288, 672, 1440}
4 The XXLTraffic Datasets
4.1 Data Collection
We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) [2]. PeMS is an online platform that collects traffic data from 19,561 sensors distributed across California state highways. These sensor locations are divided into nine districts. We downloaded all the raw data for these nine districts from the initial data release up to March 20, 2024. The system automatically generates a daily data file for each district, containing data from all sensors within each district. We have stored the complete raw data files in an open-source repository for quick access, which will be released after the publication.
Table 2: Comparison between our XXLTraffic dataset and the existing traffic datasets.
Reference Dataset Samples Nodes Time Interval Time Span Time Period
DCRNN [20] METR-LA 34,272 207 5 mins 4 months 03/2012 - 06/2012
PEMS-BAY 52,116 325 5 mins 6 months 01/2017 - 05/2017
LSTNet [14] Traffic 17,544 862 1 hour 2 years 01/2015 - 12/2016
STGCN [41] PEMSD7(M) 12,672 228 5 mins 2 months 05/2012 - 06/2012
PEMSD7(L) 12,672 1026 5 mins 2 months 05/2012 - 06/2012
ASTGCN [9] PEMSD4-I 17,002 228 5 mins 2 months 01/2018 - 02/2018
PEMSD8-I 17,856 1,979 5 mins 2 months 07/2016 - 08/2016
STSGCN [32] PEMS03 26,208 358 5 mins 11 months 01/2018 - 11/2018
PEMS04 16,992 307 5 mins 2 months 01/2018 - 02/2018
PEMS07 28,224 883 5 mins 2 months 05/2017 - 06/2017
PEMS08 17,856 170 5 mins 2 months 07/2016 - 08/2016
Large-ST [25] CA 525,888 8,600 5 mins 5 years 01/2017 - 12/2021
GLA 525,888 3,834 5 mins 5 years 01/2017 - 12/2021
GBA 525,888 2,352 5 mins 5 years 01/2017 - 12/2021
SD 525,888 716 5 mins 5 years 01/2017 - 12/2021
Ours Full_PEMS03 2,419,488 1809 5 mins 23.00 years 03/2001 - 03/2024
Full_PEMS04 2,287,872 4,089 5 mins 21.75 years 06/2002 - 03/2024
Full_PEMS05 1,998,720 573 5 mins 19.00 years 03/2005 - 03/2024
Full_PEMS06 1,945,728 705 5 mins 18.50 years 09/2005 - 03/2024
Full_PEMS07 2,287,872 4,888 5 mins 21.75 years 06/2002 - 03/2024
Full_PEMS08 2,419,488 2,059 5 mins 23.00 years 03/2001 - 03/2024
Full_PEMS10 1,998,720 1,378 5 mins 19.00 years 03/2005 - 03/2024
Full_PEMS11 2,261,376 1,440 5 mins 21.50 years 09/2002 - 03/2024
Full_PEMS12 2,331,360 2,587 5 mins 22.16 years 01/2002 - 03/2024
As illustrated in Table 2, our collected dataset significantly exceeds existing datasets in terms of both temporal coverage and the number of spatial nodes. The dataset sample will be available on: https://github.com/cruiseresearchgroup/XXLTraffic, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
4.2 Data Preprocessing
Based on the 23 years of raw data we collected, we conducted rigorous data filtering and aggregation. The PeMS system has continuously evolved, expanding from a few sensors in 2001 to over 4,000 sensors in some districts today. To support our setting of extremely long forecasting with gaps, we selected a subset of sensors that were installed in the early stages and have consistently collected new data up to the present(named gap dataset), which is shown in the Appendix. This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Figure 2. Utilizing the gap dataset, we performed both hourly and daily aggregations, which will be employed for gap-free long-term forecasting benchmarking. We will provide standard long-term forecasting benchmarks for both the hourly and daily datasets.
4.3 Data Overview
The XXLTraffic dataset is distributed across highways in the state of California, as illustrated in the Figure 3(a). The nine colors represent nine districts. From Figures 3(b),  3(c), and 3(d), we can clearly observe the evolutionary growth of the sensors. The sensors are extensively distributed across both urban and suburban areas, offering diverse modalities. Additionally, the sensors are densely interconnected, enabling the formation of a high-quality traffic graph dataset.
(b) District 08 in 2005
(c) District 08 in 2015
(d) District 08 in 2024
Figure 3: XXLTraffic dataset overview and its evolving development. This figure provides a global overview and two local overviews, showcasing the diversity of sensor distribution. The lower parts highlights a selected region to illustrate the growth and changes in traffic sensors over time.
It is evident that sensors at the same location may collect completely different distributions over the course of urban evolution. As shown in Figure 4, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution. The temporal changes causing domain shifts present a significant challenge for our extremely long forecasting.
Figure 4: Sensor traffic status distribution of District 8 from 2005 to 2024. While some sensors exhibit minimal changes, others show significant distribution differences, regardless of whether they are in low-traffic or high-traffic areas. This presents substantial challenges for extremely long forecasting with long gaps.
4.4 XXLTraffic Licence
The XXLTraffic dataset is licensed under CC BY-NC 4.0 International: https://creativecommons.org/licenses/by-nc/4.0. Our code is available under the MIT License: https://opensource.org/licenses/MIT. Please check the official repositories for the licenses of any specific baseline methods used in our codebase.
5 Experiments
We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset. Additionally, referring to the definition in Figure 2, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Figure 5.
Figure 5: Problem definition. The yellow boxes represent typical predictions, the gray boxes denote gap periods between observation and prediction, and the blue boxes indicate extended predictions.
5.1 Datasets
We conducted experiments on all proposed sub-datasets. To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, 07, and 08 (widely recognized as PEMS03/04/07/08) for our experiments, using the gap, hourly, and daily datasets from these districts. Results for other datasets are presented in Appendix. All sub-datasets were divided into training, validation, and test sets using a 6:2:2 ratio. For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results. In our experiments, due to GPU memory constraints, we pruned PEMS07_ours and PEMS12_ours by removing nodes with a missing data ratio greater than 10%, reducing the node number to 1,613 and 867. The details of the datasets used in our benchmarking is in Appendix A.1.
Table 3: Comparison in gap dataset. The bold text indicates the best.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 96 1.457 0.913 1.636 0.989 1.500 0.933 1.301 0.906
192 1.472 0.922 1.597 0.970 1.542 0.945 1.266 0.867
336 1.434 0.913 1.512 0.935 1.531 0.935 1.137 0.807
1.5-year gap 96 1.485 0.942 1.879 1.078 1.653 0.976 1.467 0.945
192 1.442 0.934 1.753 1.017 1.642 0.968 1.460 0.923
336 1.446 0.938 1.662 0.987 1.632 0.968 1.345 0.882
2-year gap 96 1.359 0.833 1.844 1.048 1.568 0.954 1.328 0.894
192 1.216 0.772 1.729 1.008 1.473 0.896 1.235 0.837
336 1.294 0.801 1.614 0.960 1.407 0.870 1.220 0.838
PEMS04_gap 1-year gap 96 1.325 0.914 1.644 1.037 1.396 0.955 0.819 0.717
192 1.438 0.642 1.587 1.018 1.440 0.965 0.941 0.767
336 1.424 0.935 1.447 0.960 1.421 0.954 0.853 0.730
1.5-year gap 96 1.204 0.890 2.014 1.193 1.488 0.995 0.981 0.779
192 0.982 0.787 1.649 1.038 1.301 0.913 0.867 0.735
336 0.961 0.792 1.352 0.929 1.298 0.908 0.762 0.679
2-year gap 96 1.220 0.893 1.652 1.074 1.446 0.977 0.909 0.755
192 1.004 0.807 1.189 0.870 1.268 0.893 0.909 0.747
336 1.198 0.864 1.584 1.016 1.269 0.898 0.898 0.736
PEMS07_gap 1-year gap 96 1.719 1.006 1.756 1.024 1.680 1.024 1.215 0.867
192 1.637 0.996 1.816 1.061 1.776 1.043 1.202 0.852
336 1.637 0.991 1.784 1.028 1.774 1.034 1.098 0.774
1.5-year gap 96 1.571 0.972 1.705 1.016 1.585 0.987 1.184 0.844
192 1.578 0.982 1.657 1.004 1.624 0.996 1.088 0.800
336 1.818 1.127 1.712 1.028 1.668 1.036 0.991 0.729
2-year gap 96 5.411 1.319 2.055 1.149 1.746 1.048 1.099 0.802
192 12.620 1.370 2.006 1.128 1.705 1.002 1.284 0.875
336 9.614 1.378 1.812 1.057 1.605 0.968 0.972 0.734
PEMS08_gap 1-year gap 96 5.411 1.319 1.514 0.950 1.771 1.058 1.153 0.843
192 12.620 1.370 1.499 0.900 1.762 1.059 1.202 0.845
336 9.614 1.378 1.654 0.984 1.961 1.131 1.184 0.849
1.5-year gap 96 4.453 1.061 2.286 1.196 1.978 1.072 1.428 0.901
192 9.413 1.046 1.713 0.971 1.606 0.928 1.314 0.833
336 10.457 1.063 1.890 1.488 1.736 0.985 1.320 0.836
2-year gap 96 5.117 0.953 2.400 1.232 1.969 1.069 1.474 0.895
192 12.769 0.933 1.974 1.070 1.670 0.942 1.505 0.883
336 13.382 0.946 1.936 1.056 1.628 0.926 1.464 0.874
5.2 Baselines
In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures. Autoformer [37], an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models. iTransformer [26] is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens. DLinear [43] challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points. It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers. Lastly, Mamba [8], a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies.
5.3 Implementation Details
We adopted the default settings of the Time-Series-Library [36] to conduct a comprehensive comparison of baselines. We use the results of five random seeds as the average. We used 96 time steps as input and 336 time steps as ground truth. The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
5.4 Results of Extremely Long Forecasting with gaps
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics to evaluate performance, averaging results across different seeds. It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps. These baseline results also indicate that traditional SOTA rankings and methodologies are no longer effective. Notably, Autoformer, which performs the worst under conventional data and settings, shows the best performance in our setting. This is understandable because Autoformer’s core technology focuses on exploring the correlation within the data itself. This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data.
Considering that our dataset provides an extensive temporal range for training, we can theoretically extend the input step length considerably. We extended the 96 steps input from Table 3 to a maximum of 1440 steps for testing. As shown in Table 4, the performance improves significantly with the increase in input step length, further demonstrating the substantial potential of our dataset for deep exploration.
Table 4: Results of ablation study with 4 different input step lengths
DLinear (Input Length) 96 192 336 720
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 96 1.500 0.933 1.455 0.912 1.462 0.906 1.392 0.887
192 1.542 0.945 1.147 0.910 1.457 0.906 1.445 0.906
336 1.531 0.935 1.447 0.907 1.462 0.906 1.439 0.904
5.5 Results of Hourly and Daily Forecasting
We believe that both the hourly and daily datasets are equally significant. Multi-scale, diverse datasets can provide the community with valuable references. We observe that the performance degrades progressively from the hourly to the daily to the gap datasets. Smaller time scales help reduce complexity and uncertainty, thereby improving prediction accuracy. Research has shown that clustering at different scales can enhance model performance [34]. Therefore, our aggregated version of the data will contribute new external features to the community.
Table 5: Comparison in hourly and daily datasets
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03_agg Hourly 96 0.144 0.222 0.530 0.535 0.159 0.222 0.241 0.346
192 0.173 0.237 0.215 0.289 0.153 0.208 0.235 0.340
336 0.158 0.220 0.519 0.527 0.167 0.216 0.260 0.362
Daily 96 0.754 0.503 0.606 0.419 0.602 0.426 0.771 0.537
192 0.968 0.604 0.781 0.500 0.794 0.509 0.897 0.577
336 1.210 0.706 0.967 0.579 0.984 0.584 1.058 0.630
PEMS04_agg Hourly 96 0.137 0.244 0.240 0.339 0.161 0.245 0.178 0.295
192 0.132 0.239 0.260 0.361 0.142 0.223 0.175 0.288
336 0.121 0.216 0.226 0.413 0.145 0.226 0.197 0.316
Daily 96 0.672 0.499 0.534 0.442 0.507 0.415 0.644 0.506
192 0.720 0.549 0.634 0.508 0.610 0.483 0.749 0.580
336 0.795 0.602 0.706 0.555 0.663 0.522 0.728 0.569
PEMS07_agg Hourly 96 0.212 0.302 0.375 0.425 0.203 0.259 0.307 0.390
192 0.201 0.288 0.297 0.368 0.182 0.231 0.313 0.391
336 0.191 0.245 0.126 0.156 0.190 0.241 0.291 0.364
Daily 96 1.719 0.736 1.426 0.613 1.414 0.606 1.703 0.762
192 2.005 0.842 1.772 0.730 1.756 0.720 1.903 0.804
336 2.290 0.949 2.078 0.819 2.051 0.813 2.171 0.884
PEMS08_agg Hourly 96 0.245 0.287 0.363 0.379 0.253 0.272 0.305 0.377
192 0.269 0.292 0.341 0.354 0.254 0.259 0.340 0.401
336 0.283 0.298 0.369 0.369 0.281 0.272 0.452 0.468
Daily 96 0.870 0.558 0.766 0.486 0.746 0.478 0.913 0.604
192 1.023 0.635 0.911 0.554 0.906 0.548 1.026 0.648
336 1.161 0.697 1.024 0.617 1.022 0.602 1.127 0.689
6 Prospects and Constraints
Prospects. Our dataset spans the longest time period among existing datasets, and it is not only the largest spatially, but also evolving in growth. It can continue to update in line with the updates from the PeMS system in the future. It is specifically designed for various complex scenarios, such as those already mentioned, including extremely long forecasting with long gaps, and hourly and daily predictions. Additionally, zero-shot forecasting designed for evolving growth scenarios will also be included in Appendix.
Constraints. The limitations of our dataset are also quite evident. Due to the sheer size of our dataset, it requires more computational resources. However, with the maturation of large language models and foundational models, we believe its large volume will become an advantage, contributing more diverse data to the spatio-temporal large model community.
Acknowledgments and Disclosure of Funding
We would like to acknowledge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program and the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005). We acknowledge the resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] We have discussed and identified no negative societal impact.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 4.1
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5.1 and Section 5.3
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] As discussed in Section 5.3, our experiments area averaged over five random seeds.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 5.3
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We have cited the creators in Section 4.1
    2. (b)
      Did you mention the license of the assets? [Yes] See Section 4.4. Our dataset complies with the CC BY-NC 4.0 license, and the code adheres to the MIT license.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [No]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Please refer to Section 4.1 for the detailed discussion.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No] This dataset is distributed across the highway, collecting traffic speed, traffic flow, and traffic occupancy. It does not contain personally identifiable information or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Appendix
A.1 Dataset
To support our setting of extremely long forecasting with gaps, we selected a subset installed in the early stages and have consistently collected new data up to the present, which is shown as follows:
Table 6: Overview of Gap, Hourly, and Daily Aggregated Data Based on Processed Raw Data
Datasets(Gap/Hour/Day) Time Period Nodes
PEMS03_gap&agg 03/2001 - 03/2024 151
PEMS04_gap&agg 06/2002 - 03/2024 822
PEMS05_gap&agg 03/2012 - 03/2024 103
PEMS06_gap&agg 12/2009 - 03/2024 130
PEMS07_gap&agg 06/2002 - 03/2024 3062
PEMS08_gap&agg 03/2001 - 03/2024 212
PEMS10_gap&agg 06/2007 - 03/2024 107
PEMS11_gap&agg 09/2002 - 03/2024 521
PEMS12_gap&agg 01/2002 - 03/2024 1543
A.2 Results
Here we provided the results of District 5,6,10,11,12, as shown in Table 7 and Table 8:
Table 7: Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS05_agg Hourly 96 0.121 0.205 0.226 0.324 0.148 0.236 0.155 0.272
192 0.118 0.197 0.214 0.312 0.132 0.216 0.159 0.268
336 0.115 0.194 0.220 0.318 0.134 0.217 0.167 0.274
Daily 96 0.655 0.511 0.654 0.510 0.607 0.477 0.650 0.522
192 0.798 0.598 0.775 0.576 0.698 0.535 0.643 0.512
336 0.907 0.614 0.787 0.582 0.745 0.556 0.764 0.578
PEMS06_agg Hourly 96 0.142 0.234 0.269 0.336 0.188 0.254 0.164 0.271
192 0.138 0.218 0.256 0.326 0.166 0.227 0.188 0.291
336 0.137 0.207 0.226 0.413 0.171 0.227 0.197 0.316
Daily 96 0.516 0.419 0.414 0.389 0.405 0.340 0.518 0.437
192 0.642 0.485 0.543 0.468 0.510 0.395 0.580 0.460
336 0.734 0.519 0.675 0.492 0.596 0.437 0.658 0.481
PEMS10_agg Hourly 96 0.213 0.256 0.391 0.413 0.272 0.309 0.260 0.346
192 0.205 0.250 0.387 0.412 0.246 0.277 0.380 0.417
336 0.211 0.255 0.394 0.413 0.258 0.281 0.329 0.386
Daily 96 1.161 0.671 0.926 0.513 0.951 0.567 1.079 0.647
192 1.459 0.784 1.414 0.730 1.228 0.681 1.429 0.786
336 1.715 0.855 1.478 0.768 1.451 0.751 1.552 0.817
PEMS11_agg Hourly 96 - 0.472 - 0.538 - 0.306 - 0.800
192 - 0.470 - 0.443 - 0.291 - 0.742
336 - 0.487 - 0.435 - 0.300 - 0.745
Daily 96 - - - - - - - -
192 - - - - - - - -
336 - - - - - - - -
PEMS12_agg Hourly 96 0.145 0.237 0.083 0.157 0.174 0.242 0.188 0.287
192 0.142 0.226 0.091 0.159 0.154 0.216 0.196 0.289
336 0.146 0.217 0.104 0.170 0.162 0.219 0.209 0.298
Daily 96 1.373 0.621 1.456 0.622 1.052 0.510 1.348 0.649
192 1.722 0.726 1.675 0.678 1.403 0.611 1.548 0.686
336 2.066 0.823 1.984 0.641 1.654 0.675 1.801 0.746
Table 8: Comparison in gap dataset in District 05,06,10,11,12. The bold text indicates the best.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS05_gap 1-year gap 96 2.079 1.209 1.945 1.164 1.291 0.916 1.065 0.796
192 2.132 1.256 1.984 1.185 1.750 1.099 1.063 0.809
336 2.377 1.340 2.067 1.234 1.894 1.144 1.135 0.827
1.5-year gap 96 1.852 1.122 1.879 1.078 1.683 1.054 1.060 0.785
192 1.929 1.182 1.593 1.032 1.633 1.045 0.912 0.712
336 2.370 1.313 2.214 1.071 0.794 1.184 1.345 0.882
2-year gap 96 1.868 1.106 1.580 0.969 1.602 1.018 0.828 0.672
192 2.219 1.274 1.481 0.958 1.589 1.027 1.018 0.772
336 2.695 1.212 2.207 1.201 1.922 1.139 1.186 0.839
PEMS06_gap 1-year gap 96 1.806 1.066 0.875 0.692 1.173 0.837 1.216 0.859
192 1.928 1.112 1.227 0.848 1.410 0.942 0.961 0.751
336 2.181 1.212 1.594 1.003 1.501 0.976 0.992 0.769
1.5-year gap 96 1.549 0.997 1.331 0.891 1.484 0.963 0.885 0.710
192 1.746 1.054 1.077 0.778 1.353 0.920 1.010 0.768
336 1.605 1.018 1.500 0.961 1.587 1.011 0.955 0.739
2-year gap 96 1.226 0.851 1.864 1.106 1.691 1.033 1.013 0.768
192 0.949 0.720 1.343 0.879 1.259 0.858 0.853 0.691
336 0.945 0.710 1.550 0.970 1.415 0.934 0.955 0.739
PEMS10_gap 1-year gap 96 2.310 1.233 0.882 0.698 1.466 0.963 1.208 0.865
192 2.878 1.396 1.606 0.993 1.834 1.106 1.152 0.860
336 2.584 1.327 1.413 0.825 1.887 1.121 1.130 0.830
1.5-year gap 96 2.181 1.206 1.812 1.052 1.765 1.057 1.368 0.913
192 2.525 1.318 1.645 0.986 1.791 1.069 1.151 0.828
336 2.488 1.310 2.051 1.134 1.976 1.139 1.111 0.809
2-year gap 96 1.165 0.822 2.772 1.383 1.977 1.119 1.188 0.845
192 1.082 0.792 1.974 1.084 1.490 0.932 0.971 0.751
336 1.021 0.764 2.157 1.168 1.619 0.989 1.096 0.809
PEMS11_gap 1-year gap 96 5.199 0.891 5.417 0.997 5.276 0.936 4.930 0.854
192 5.300 0.920 5.504 1.029 5.393 0.974 5.114 0.867
336 5.251 0.901 5.410 0.985 5.364 0.955 4.830 0.849
1.5-year gap 96 5.871 0.968 6.045 1.296 5.914 1.224 5.691 0.878
192 5.968 1.012 6.121 1.314 5.993 1.229 5.792 0.884
336 6.167 1.074 6.214 1.326 6.025 1.299 5.947 0.967
2-year gap 96 5.914 0.996 6.136 1.318 5.945 1.243 5.761 0.978
192 6.541 1.043 6.213 1.327 6.014 1.289 6.245 1.001
336 6.541 1.086 6.221 1.332 6.024 1.300 6.268 1.024
PEMS12_gap 1-year gap 96 1.751 1.025 1.624 1.002 1.611 1.005 1.060 0.789
192 1.726 1.024 1.424 0.929 1.537 0.972 1.150 0.840
336 1.751 1.029 1.672 1.015 1.683 1.017 0.889 0.719
1.5-year gap 96 1.554 0.967 1.479 0.921 1.468 0.910 0.954 0.875
192 1.401 0.898 1.314 0.867 1.301 0.862 0.943 0.869
336 1.417 0.906 1.322 0.869 1.298 0.859 0.921 0.846
2-year gap 96 0.956 0.704 0.876 0.671 0.872 0.664 0.846 0.659
192 0.846 0.656 0.813 0.653 0.806 0.649 0.785 0.628
336 0.814 0.628 0.789 0.631 0.776 0.624 0.754 0.617
Additionally, we provided the results of naive zero-shot forecasting, as shown in Table 9. The poor performance of this method indicates significant potential for improvement.
Table 9: Results of zero shot forecasting in PEMS03. In this study, we utilized the inputs from 151 existing nodes in PEMS03 to predict 52 new nodes, resulting in a test set comprising a total of 203 nodes.
Datasets 96 192 336
Metrics MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 4.658 3.465 4.945 3.648 5.198 3.892
1.5-year gap 4.891 3.657 5.124 3.842 5.263 3.996
2-year gap 4.547 3.410 4.895 3.539 5.103 3.758